What is x/x+2 = 5/7 and how do you determine the answer?
Found 2 solutions by stanbon, rothauserc:
Answer by stanbon(75887)

x/x+2 = 5/7
7x = 5x+10
2x =10
x = 5
==============
Cheers,
Stan H.
==========

Answer by rothauserc(4718)

when two fractions are equal we can cross multiply - that is, numerator of the first fraction times the denominator of the second fraction equals the denominator of the first fraction times the numerator of the second fraction
================================================================================
x/x+2 = 5/7, cross multiplying we get
7x = 5*(x+2)
7x = 5x +10
subtract 5x from both sides of the =
2x = 10
x = 5